Grade 12Physical Chemistry

Convert the following into basic units: (i) 28.7 pm (ii) 15.15 pm (iii) 25365 mg

sudhanshu
12 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
(i)pm=pico meter
1pm=10-12meter(m)
so 28.7 pm=28.7*10-12meter=2.87*10-11m
(ii)
15.15pm
similar way as in (i)
(iii)1mg=1 mili gram=10-3 gram(g)
so 25365mg=25365*10-3g=25.365g